Calculate the molar concentration of hydrogen ions in solution: (a) pH = 2.8 and (b) pOH = 4.2?


1
Expert's answer
2021-01-12T07:04:57-0500

[H+] is the molar hydrogen ion concentration (in mol L-1 or M).


Solution:

(a):

pH is defined as the negative logarithm of the hydrogen ion concentration:

pH = -log[H+]

Hence, [H+] can be calculated using the above equation:

[H+] = 10-pH

[H+] = 10-2.8 = 0.001585 M = 1.6×10-3 M

[H+] = 1.6×10-3 M


(b):

For any aqueous solution at 25C:

pH + pOH = 14

Hence,

pH = 14 - pOH = 14 - 4.2 = 9.8

pH is defined as the negative logarithm of the hydrogen ion concentration:

pH = -log[H+]

Hence, [H+] can be calculated using the above equation:

[H+] = 10-pH

[H+] = 10-9.8 = 1.585×10-10 M = 1.6×10-10 M

[H+] = 1.6×10-10 M


Answer:

(a): [H+] = 1.6×10-3 M

(b): [H+] = 1.6×10-10 M